# Test-Time Training on Video Streams

### v1

@@ 1 Introduction
Most models in machine learning today are fixed during deployment. As a consequence, a trained model must prepare to be robust to all possible futures. This can be hard because being ready for all futures limits the model’s capacity to be good at any particular one. But only one future actually happens at each moment. The basic idea of test-time training is to continue training on the future once it arrives in the form of a test instance [19, 61]. Since each test instance is observed without a ground truth label, training is performed with self-supervision.
@@
Our empirical evidence supports locality. The best performance is achieved through online TTT: For each $x_{t}$, only train on itself, and a small window of less than two seconds of frames immediately before $t$. We call this window of frames sliding with $t$ the explicit memory. The optimal explicit memory needs to be short term – in the language of continual learning, some amount of forgetting is actually beneficial. This contrasts with findings in the conventional setting of continual learning [37, 42, 33], but is consistent with recent studies in neuroscience [22].
@@
Experiments in this paper are also of practical interests, besides conceptual ones. Models for many computer vision tasks are trained with large datasets of still images, e.g. COCO [38] for segmentation, but deployed on video streams. The default is to naively run such models frame-by-frame, since averaging across a sliding window of predictions, a.k.a. temporal smoothing, offers little improvement. Online TTT significantly improves prediction quality on three real-world video datasets, for four tasks: semantic, instance and panoptic segmentation, and colorization. Figure 2 visualizes results for the first three tasks with reliable metrics.
@@
One of the most popular forms of self-supervision in computer vision is image reconstruction: removing parts of the input image, then predicting the removed content [68, 48, 6, 75]. Recently, a class of deep learning models called masked autoencoders (MAE) [27], using reconstruction as the self-supervised task, has been highly influential. TTT-MAE [19] adopts these models for test-time training using reconstruction. The main task in [19] is object recognition. Inspired by the empirical success of TTT-MAE, we make it the inner loop of online TTT, and extend [19] to other main tasks such as segmentation.
Prior work [61] experiments with online TTT (without explicit memory) in the streaming setting, but each $x_{t}$ is drawn independently from the same test distribution. This test distribution is created by adding some synthetic corruption, e.g. Gaussian noise, to a test set of still images, e.g. ImageNet test set [28]. Therefore, all $x_{t}$s belong to the same “future”, and locality is meaningless: TTT on as many $x_{t}$s as possible achieves the best performance by learning to ignore the corruption. TTT on actual video streams is fundamentally different and much more natural.
More recently, [69] also experiments in the streaming setting. While the $x_{t}$s here are not independent, these videos are short clips, again simulated to contain synthetic corruptions, e.g. CityScapes with Artificial Weather. Therefore, like in [61], each corruption moves all $x_{t}$s into almost the same “future”, which they call a domain. Since performance drop is caused by that shared corruption, it is best recovered by training on all $x_{t}$s. Their only dataset without corruptions (CityScapes) sees little improvement (1.4% relative). There is no mentioning of locality, our basic concept of interest.
@@ 2.1 Continual Learning
In the field of continual a.k.a. lifelong learning, a model learns a sequence of tasks in temporal order, and is asked to perform well on all of them [65, 23]. Here is the conventional setting: Each task is defined by a data distribution $P_{t}$, which produces a training set $D^{\text{tr}}_{t}$ and a test set $D^{\text{te}}_{t}$. At each time $t$, the model is evaluated on all the test sets $D^{\text{te}}_{1},\dots,D^{\text{te}}_{t}$ of the past and present, and average performance is reported.
The basic solution is to simply train the model on all of $D^{\text{tr}}_{1},\dots,D^{\text{tr}}_{t}$, which collectively have the same distribution as all the test sets. This is often referred to as the oracle with infinite memory (a.k.a. replay buffer) that remembers everything. However, due to memory constraints, the model at time $t$ is only allowed to train on $D^{\text{tr}}_{t}$. More advanced solutions, therefore, focus on how to retain memory of past data only with model parameters [53, 37, 42, 56, 33, 20].
Some of the literature extends beyond the conventional setting. [1] uses continuous instead of discrete tasks across time. [50] and [17] perform self-supervised learning on unlabeled training sets, and evaluate the learned features on the test sets. [29], [36] and [47] use a labeled training set $D^{\text{tr}}_{0}$ in addition to unlabeled training sets $D^{\text{tr}}_{1},\dots,D^{\text{tr}}_{t}$, connecting with unsupervised domain adaptation. [15] uses alternative metrics, e.g. forward transfer, to justify forgetting for reasons other than computational.
Much of continual learning is motivated by the hope to understand human memory and generalization through the lens of artificial intelligence [26, 13]. Our work shares the same motivation, but focuses on test-time training, without distinct splits of training and test sets.
@@ 2.2 Test-Time Training
One of the earliest algorithms for training at test time is Bottou and Vapnik [7]: For each test input, train on its neighbors before making a prediction. Their paper, titled Local Learning, is the first to articulate locality as a basic concept for parametric models in machine learning. This continues to be effective for support vector machines (SVM) [79] and large language models [25]. Another line of work called transductive learning uses test data to add constraints to the margin of SVMs [31, 11, 66]. The principle of transduction, as stated by Vapnik, also emphasizes locality  [18, 67]: "Try to get the answer that you really need but not a more general one."
In computer vision, the idea of training at test time has been well explored for specific applications [30, 57, 46, 73], especially depth estimation [62, 63, 82, 84, 43]. Our paper extends TTT-MAE [19], detailed in Section 3. TTT-MAE, in turn, is inspired by the work Sun et al. [61], which proposed the general framework for test-time training with self-supervision, regardless of application. The particular self-supervised task used in [61] is rotation prediction [21]. Many other papers have followed this framework since then [24, 60, 40, 77], including [69] on videos discussed in Section 1, and [5] which we discuss next.
In [5], each video is treated as a dataset of unordered frames instead of a stream. In particular, there is no concept of past vs. future frames. The same model is used on the entire video. In contrast, our paper emphasizes locality. We have access to only the current and past frames, and our model keeps learning over time. In addition, all of our results are on real world videos, while [5] experiment on videos with artificial corruptions. These corruptions are also i.i.d. across frames.
Our paper is very much inspired by [45]. To make video segmentation more efficient, [45] makes predictions frame-by-frame using a small student model. If the student is not confident, it queries an expensive teacher model, and then trains the student to fit the prediction from the teacher online. Thanks to temporal smoothness, the student can generalize confidently across many frames without querying the teacher, so learning and predicting combined is still faster than naively using the teacher at every frame. Our method only consists of one model, which learns from a self-supervised task instead of a teacher model. Rather than focusing on computational efficiency as in [45], the main goal of our paper is to improve inference quality. Behind their particular algorithm, however, we see the shared idea of locality, regardless of the form of supervision.
@@ 3 Background: TTT-MAE
Our paper extends the work of Test-Time Training with Masked Autoencoders (TTT-MAE) [19], and uses TTT-MAE as the inner loop when updating the model for each frame. This section briefly describes TTT-MAE, as background for our extension. Figure 3 illustrates the process of TTT-MAE.
The architecture for TTT with self-supervision [61] is Y-shaped with a stem and two heads: a prediction head $g$ for the self-supervised task, a prediction head $h$ for the main task, and a feature extractor $f$ as the stem. The output features of $f$ are shared between $g$ and $h$ as input. For TTT-MAE, the self-supervised task is masked image reconstruction [27]. Following standard terminology for autoencoders, $f$ is also called the encoder, and $g$ the decoder.
@@ 3.1 Training-Time Training
There are three widely accepted ways to optimize the model components ($f$, $g$, $h$) at training time: joint training, probing, and fine-tuning. Fine-tuning is unsuitable for TTT, because it makes $h$ rely too much on features that are used by the main task. Our paper uses joint training, described in Section 4. In contrast, [19] uses probing, which we describe next for completeness.
To prepare for probing, the common practice is to first train $f$ and $g$ with $\ell_{s}$ on the training set without ground truth. This preparation stage is also called self-supervised pre-training. TTT-MAE [19] uses the encoder and decoder already pre-trained by [27], denoted by $f_{0}$ and $g_{0}$. During probing, the main task head $h$ is then trained separately by optimizing for $\ell_{m}(h\circ f_{0}(x),y)$, on the training set with ground truth. $f_{0}$ is kept frozen. We denote $h_{0}$ as the main task head after probing. Since $h_{0}$ has been trained for the main task using features from $f_{0}$ as input, $h_{0}\circ f_{0}$ can be directly applied on each test image as a baseline without TTT, keeping the parameters of $f_{0}$ and $h_{0}$ fixed.
Figure 3: Training a masked autoencoder (MAE) to reconstruct each test image at test time. Reconstructed images on the right visualize the progress of gradient descent on this one-sample learning problem. For each test image, TTT-MAE [19] first masks out majority of the patches. The masked image is given as input to the autoencoder, which then reconstructs those masked patches. The reconstruction loss is the pixel-wise mean squared error between the original and reconstructed patches. Loss on the main task – panoptic segmentation – also falls as reconstruction gets better. The unmasked patches are not shown on the right since they are not part of the reconstruction loss.
@@ 3.2 Test-Time Training
then makes the final prediction $h_{0}\circ f^{\prime}(x^{\prime})$. Crucially, the gradient-based optimization process always starts from $f_{0}$ and $g_{0}$. When evaluating on a test set, TTT-MAE [19] always discards $f^{\prime}$ and $g^{\prime}$ after making a prediction on each test input $x^{\prime}$, and resets the weights back to $f_{0}$ and $g_{0}$ for the next test input. By test-time training on the test inputs independently, we do not assume that they can help each other.
In the original MAE design [27], $g$ is very small relative to $f$, and only the visible patches, e.g. 20%, are processed by $f$. Therefore the overall computational cost of training for the self-supervised task in only a fraction, e.g. $25\%$, of training for the main task. In addition to speeding up training-time training for reconstruction, this reduces the extra test-time cost of TTT-MAE. Each gradient step at test time, counting both forward and backward, costs only half the time of forward prediction for the main task.
@@ 4 Test-Time Training on Video Streams
Now we describe our algorithm for this streaming setting. At a high level, our algorithm simply amounts to an outer loop wrapped around TTT-MAE [19]. In practice, making it work involves many design choices.
@@ 4.1 Training-Time Training
At training time, our algorithm jointly optimize all three model components in a single stage, end-to-end, on both the self-supervised task and main task. This is called joint training. While joint training was also an option for prior work on TTT-MAE [19], empirical experience at the time indicated that probing (see Section 3) performed better. In this paper, however, we successfully tune joint training to be as effective as probing, and therefore default to joint training because it is simpler than the two-stage process of probing.
@@
The summation is over the training set with $n$ samples, each consisting of input $x_{i}$ and label $y_{i}$. As discussed in Section 3, $\tilde{x}_{i}$ is $x_{i}$ transformed as input for the self-supervised task. In the case of MAE, $\tilde{x}_{i}$ is obtained by masking $80\%$ of the patches in $x_{i}$. Note that although the test instances come from video streams, training-time training uses labeled, still images, e.g. in the COCO training set [38], instead of unlabeled videos.
@@ 4.2 Test-Time Training
Implicit memory. The most natural improvement is to simply not reset the model parameters between timesteps. That is, to initialize test-time training at timestep $t$ with $f_{t-1}$ and $g_{t-1}$, instead of $f_{0}$ and $g_{0}$. This creates an implicit memory, since information carries over from the previous parameters, optimized on previous frames. It also happens to be more biologically plausible: we humans do not constantly reset our minds. In prior work [61], TTT with implicit memory is called the “online” version, in contrast to the “standard” version with reset, for the setting of independent images without temporal smoothness.
@@ 4.3 Implementation Details
In principle, our method is applicable to any architecture. We use Mask2Former [10], which has achieved state-of-the-art performance on many semantic, instance and panoptic segmentation benchmarks. Our Mask2Former uses a Swin-S [41] backbone – in our case, this is also the shared encoder $f$. Everything following the backbone in the original architecture is taken as the main task head $h$, and our decoder $g$ copies the architecture of $h$ except the last layer that maps into pixel space for reconstruction. Joint training starts from their model checkpoint, which has already been trained for the main task. Only $g$ is initialized from scratch.
Following [27], we split each input into patches, and mask out 80% of them. However, unlike the Vision Transformers [16] used in [27], Swin Transformers use convolutions. Therefore, we must take the entire image as input (with the masked patches in black) instead of only the unmasked patches. Following [48], we use a fourth channel of binaries to indicate if the corresponding input pixels are masked. The model parameters for the fourth channel are initialized from scratch before joint training. If a completely transformer-based architecture for segmentation becomes available in the future, our method would like become even faster, by not encoding the masked patches [27, 19].
@@
Independent frames Main Task Only [10] 13.4 14.6 53.8 52.5 1.8
@@
TTT on video streams LN Adapt [54] 13.3 15.3 53.8 52.5 2.0
Tent [70] 13.3 15.2 53.8 52.2 2.8
Tent Class [70] 13.5 15.3 53.8 52.5 3.7
Self-Train [69] - - 54.7 54.0 6.6
Self-Train Class [69] - - 54.1 53.6 6.9

### v3

@@ 1 Introduction
Most models in machine learning today are fixed during deployment. As a consequence, a trained model must prepare to be robust to all possible futures. This can be hard because being ready for all futures limits the model’s capacity to be good at any particular one. But only one future actually happens at each moment. The basic idea of test-time training is to continue training on the future once it arrives in the form of a test instance [ttt-mae, sun2020test]. Since each test instance is observed without a ground truth label, training is performed with self-supervision.
@@
Our empirical evidence supports locality. The best performance is achieved through online TTT: For each $x_{t}$, only train on itself, and a small window of less than two seconds of frames immediately before $t$. We call this window of frames sliding with $t$ the explicit memory. The optimal explicit memory needs to be short term – in the language of continual learning, some amount of forgetting is actually beneficial. This contrasts with findings in the conventional setting of continual learning [li2017learning, lopez2017gradient, kirkpatrick2017overcoming], but is consistent with recent studies in neuroscience [gravitz2019importance].
@@
Experiments in this paper are also of practical interests, besides conceptual ones. Models for many computer vision tasks are trained with large datasets of still images, e.g. COCO [lin2014microsoft] for segmentation, but deployed on video streams. The default is to naively run such models frame-by-frame, since averaging across a sliding window of predictions, a.k.a. temporal smoothing, offers little improvement. Online TTT significantly improves prediction quality on three real-world video datasets, for four tasks: semantic, instance and panoptic segmentation, and colorization. Figure 2 visualizes results for the first three tasks with reliable metrics.
@@
One of the most popular forms of self-supervision in computer vision is image reconstruction: removing parts of the input image, then predicting the removed content [denoisingautoencoder, pathak2016context, beit, simmim]. Recently, a class of deep learning models called masked autoencoders (MAE) [mae], using reconstruction as the self-supervised task, has been highly influential. TTT-MAE [ttt-mae] adopts these models for test-time training using reconstruction. The main task in [ttt-mae] is object recognition. Inspired by the empirical success of TTT-MAE, we make it the inner loop of online TTT, and extend [ttt-mae] to other main tasks such as segmentation.
Prior work [sun2020test] experiments with online TTT (without explicit memory) in the streaming setting, but each $x_{t}$ is drawn independently from the same test distribution. This test distribution is created by adding some synthetic corruption, e.g. Gaussian noise, to a test set of still images, e.g. ImageNet test set [imagenet_c]. Therefore, all $x_{t}$s belong to the same “future”, and locality is meaningless: TTT on as many $x_{t}$s as possible achieves the best performance by learning to ignore the corruption. TTT on actual video streams is fundamentally different and much more natural.
More recently, [volpi2022road] also experiments in the streaming setting. While the $x_{t}$s here are not independent, these videos are short clips, again simulated to contain synthetic corruptions, e.g. CityScapes with Artificial Weather. Therefore, like in [sun2020test], each corruption moves all $x_{t}$s into almost the same “future”, which they call a domain. Since performance drop is caused by that shared corruption, it is best recovered by training on all $x_{t}$s. Their only dataset without corruptions (CityScapes) sees little improvement (1.4% relative). There is no mentioning of locality, our basic concept of interest.
@@ 2.1 Continual Learning
In the field of continual a.k.a. lifelong learning, a model learns a sequence of tasks in temporal order, and is asked to perform well on all of them [van2019three, hadsell2020embracing]. Here is the conventional setting: Each task is defined by a data distribution $P_{t}$, which produces a training set $D^{\text{tr}}_{t}$ and a test set $D^{\text{te}}_{t}$. At each time $t$, the model is evaluated on all the test sets $D^{\text{te}}_{1},\dots,D^{\text{te}}_{t}$ of the past and present, and average performance is reported.
The basic solution is to simply train the model on all of $D^{\text{tr}}_{1},\dots,D^{\text{tr}}_{t}$, which collectively have the same distribution as all the test sets. This is often referred to as the oracle with infinite memory (a.k.a. replay buffer) that remembers everything. However, due to memory constraints, the model at time $t$ is only allowed to train on $D^{\text{tr}}_{t}$. More advanced solutions, therefore, focus on how to retain memory of past data only with model parameters [santoro2016meta, li2017learning, lopez2017gradient, shin2017continual, kirkpatrick2017overcoming, gidaris2018dynamic].
Some of the literature extends beyond the conventional setting. [aljundi2019task] uses continuous instead of discrete tasks across time. [purushwalkam2022challenges] and [fini2022self] perform self-supervised learning on unlabeled training sets, and evaluate the learned features on the test sets. [hoffman2014continuous], [li2020online] and [panagiotakopoulos2022online] use a labeled training set $D^{\text{tr}}_{0}$ in addition to unlabeled training sets $D^{\text{tr}}_{1},\dots,D^{\text{tr}}_{t}$, connecting with unsupervised domain adaptation. [diaz2018don] uses alternative metrics, e.g. forward transfer, to justify forgetting for reasons other than computational.
Much of continual learning is motivated by the hope to understand human memory and generalization through the lens of artificial intelligence [hassabis2017neuroscience, de2021continual]. Our work shares the same motivation, but focuses on test-time training, without distinct splits of training and test sets.
@@ 2.2 Test-Time Training
One of the earliest algorithms for training at test time is Bottou and Vapnik [bottou1992local]: For each test input, train on its neighbors before making a prediction. Their paper, titled Local Learning, is the first to articulate locality as a basic concept for parametric models in machine learning. This continues to be effective for support vector machines (SVM) [zhang2006svm] and large language models [hardt2023test]. Another line of work called transductive learning uses test data to add constraints to the margin of SVMs [joachims2002learning, collobert2006large, vapnik2013nature]. The principle of transduction, as stated by Vapnik, also emphasizes locality  [Gammerman98learningby, vapnik_book]: "Try to get the answer that you really need but not a more general one."
In computer vision, the idea of training at test time has been well explored for specific applications [jain2011online, shocher2018zero, nitzan2022mystyle, xie2023sepico], especially depth estimation [tonioni2019learning, tonioni2019real, zhang2020online, zhong2018open, luo2020consistent]. Our paper extends TTT-MAE [ttt-mae], detailed in Section 3. TTT-MAE, in turn, is inspired by the work Sun et al. [sun2020test], which proposed the general framework for test-time training with self-supervision, regardless of application. The particular self-supervised task used in [sun2020test] is rotation prediction [gidaris2018unsupervised]. Many other papers have followed this framework since then [hansen2020self, sun2021online, liu2021ttt++, yuan2023robust], including [volpi2022road] on videos discussed in Section 1, and [azimi2022self] which we discuss next.
In [azimi2022self], each video is treated as a dataset of unordered frames instead of a stream. In particular, there is no concept of past vs. future frames. The same model is used on the entire video. In contrast, our paper emphasizes locality. We have access to only the current and past frames, and our model keeps learning over time. In addition, all of our results are on real world videos, while [azimi2022self] experiment on videos with artificial corruptions. These corruptions are also i.i.d. across frames.
Our paper is very much inspired by [mullapudi2018online]. To make video segmentation more efficient, [mullapudi2018online] makes predictions frame-by-frame using a small student model. If the student is not confident, it queries an expensive teacher model, and then trains the student to fit the prediction from the teacher online. Thanks to temporal smoothness, the student can generalize confidently across many frames without querying the teacher, so learning and predicting combined is still faster than naively using the teacher at every frame. Our method only consists of one model, which learns from a self-supervised task instead of a teacher model. Rather than focusing on computational efficiency as in [mullapudi2018online], the main goal of our paper is to improve inference quality. Behind their particular algorithm, however, we see the shared idea of locality, regardless of the form of supervision.
@@ 3 Background: TTT-MAE
Our paper extends the work of Test-Time Training with Masked Autoencoders (TTT-MAE) [ttt-mae], and uses TTT-MAE as the inner loop when updating the model for each frame. This section briefly describes TTT-MAE, as background for our extension. Figure 3 illustrates the process of TTT-MAE.
The architecture for TTT with self-supervision [sun2020test] is Y-shaped with a stem and two heads: a prediction head $g$ for the self-supervised task, a prediction head $h$ for the main task, and a feature extractor $f$ as the stem. The output features of $f$ are shared between $g$ and $h$ as input. For TTT-MAE, the self-supervised task is masked image reconstruction [mae]. Following standard terminology for autoencoders, $f$ is also called the encoder, and $g$ the decoder.
@@ 3.1 Training-Time Training
There are three widely accepted ways to optimize the model components ($f$, $g$, $h$) at training time: joint training, probing, and fine-tuning. Fine-tuning is unsuitable for TTT, because it makes $h$ rely too much on features that are used by the main task. Our paper uses joint training, described in Section 4. In contrast, [ttt-mae] uses probing, which we describe next for completeness.
To prepare for probing, the common practice is to first train $f$ and $g$ with $\ell_{s}$ on the training set without ground truth. This preparation stage is also called self-supervised pre-training. TTT-MAE [ttt-mae] uses the encoder and decoder already pre-trained by [mae], denoted by $f_{0}$ and $g_{0}$. During probing, the main task head $h$ is then trained separately by optimizing for $\ell_{m}(h\circ f_{0}(x),y)$, on the training set with ground truth. $f_{0}$ is kept frozen. We denote $h_{0}$ as the main task head after probing. Since $h_{0}$ has been trained for the main task using features from $f_{0}$ as input, $h_{0}\circ f_{0}$ can be directly applied on each test image as a baseline without TTT, keeping the parameters of $f_{0}$ and $h_{0}$ fixed.
Figure 3: Training a masked autoencoder (MAE) to reconstruct each test image at test time. Reconstructed images on the right visualize the progress of gradient descent on this one-sample learning problem. For each test image, TTT-MAE [ttt-mae] first masks out majority of the patches. The masked image is given as input to the autoencoder, which then reconstructs those masked patches. The reconstruction loss is the pixel-wise mean squared error between the original and reconstructed patches. Loss on the main task – panoptic segmentation – also falls as reconstruction gets better. The unmasked patches are not shown on the right since they are not part of the reconstruction loss.
@@ 3.2 Test-Time Training
then makes the final prediction $h_{0}\circ f^{\prime}(x^{\prime})$. Crucially, the gradient-based optimization process always starts from $f_{0}$ and $g_{0}$. When evaluating on a test set, TTT-MAE [ttt-mae] always discards $f^{\prime}$ and $g^{\prime}$ after making a prediction on each test input $x^{\prime}$, and resets the weights back to $f_{0}$ and $g_{0}$ for the next test input. By test-time training on the test inputs independently, we do not assume that they can help each other.
In the original MAE design [mae], $g$ is very small relative to $f$, and only the visible patches, e.g. 20%, are processed by $f$. Therefore the overall computational cost of training for the self-supervised task in only a fraction, e.g. $25\%$, of training for the main task. In addition to speeding up training-time training for reconstruction, this reduces the extra test-time cost of TTT-MAE. Each gradient step at test time, counting both forward and backward, costs only half the time of forward prediction for the main task.
@@ 4 Test-Time Training on Video Streams
Now we describe our algorithm for this streaming setting. At a high level, our algorithm simply amounts to an outer loop wrapped around TTT-MAE [ttt-mae]. In practice, making it work involves many design choices.
@@ 4.1 Training-Time Training
At training time, our algorithm jointly optimize all three model components in a single stage, end-to-end, on both the self-supervised task and main task. This is called joint training. While joint training was also an option for prior work on TTT-MAE [ttt-mae], empirical experience at the time indicated that probing (see Section 3) performed better. In this paper, however, we successfully tune joint training to be as effective as probing, and therefore default to joint training because it is simpler than the two-stage process of probing.
@@
The summation is over the training set with $n$ samples, each consisting of input $x_{i}$ and label $y_{i}$. As discussed in Section 3, $\tilde{x}_{i}$ is $x_{i}$ transformed as input for the self-supervised task. In the case of MAE, $\tilde{x}_{i}$ is obtained by masking $80\%$ of the patches in $x_{i}$. Note that although the test instances come from video streams, training-time training uses labeled, still images, e.g. in the COCO training set [lin2014microsoft], instead of unlabeled videos.
@@ 4.2 Test-Time Training
Implicit memory. The most natural improvement is to simply not reset the model parameters between timesteps. That is, to initialize test-time training at timestep $t$ with $f_{t-1}$ and $g_{t-1}$, instead of $f_{0}$ and $g_{0}$. This creates an implicit memory, since information carries over from the previous parameters, optimized on previous frames. It also happens to be more biologically plausible: we humans do not constantly reset our minds. In prior work [sun2020test], TTT with implicit memory is called the “online” version, in contrast to the “standard” version with reset, for the setting of independent images without temporal smoothness.
@@ 4.3 Implementation Details
In principle, our method is applicable to any architecture. We use Mask2Former [cheng2021masked], which has achieved state-of-the-art performance on many semantic, instance and panoptic segmentation benchmarks. Our Mask2Former uses a Swin-S [liu2021swin] backbone – in our case, this is also the shared encoder $f$. Everything following the backbone in the original architecture is taken as the main task head $h$, and our decoder $g$ copies the architecture of $h$ except the last layer that maps into pixel space for reconstruction. Joint training starts from their model checkpoint, which has already been trained for the main task. Only $g$ is initialized from scratch.
Following [mae], we split each input into patches, and mask out 80% of them. However, unlike the Vision Transformers [dosovitskiy2020image] used in [mae], Swin Transformers use convolutions. Therefore, we must take the entire image as input (with the masked patches in black) instead of only the unmasked patches. Following [pathak2016context], we use a fourth channel of binaries to indicate if the corresponding input pixels are masked. The model parameters for the fourth channel are initialized from scratch before joint training. If a completely transformer-based architecture for segmentation becomes available in the future, our method would like become even faster, by not encoding the masked patches [mae, ttt-mae].
@@
Independent frames Main Task Only [cheng2021masked] 13.4 14.6 53.8 52.5 1.8
@@
TTT on video streams LN Adapt [schneider2020improving] 13.3 15.3 53.8 52.5 2.0
Tent [wang2020tent] 13.3 15.2 53.8 52.2 2.8
Tent Class [wang2020tent] 13.5 15.3 53.8 52.5 3.7
Self-Train [volpi2022road] - - 54.7 54.0 6.6
Self-Train Class [volpi2022road] - - 54.1 53.6 6.9